OpenStudy (anonymous):

can someone show me how to work out x^2-9 over x^2 times x^2-3x over x^2+x-12

OpenStudy (saifoo.khan):

\[\frac{x^2-9}{x^2} \times \frac{x^2-3x}{(x-3) (x+4) }\]

OpenStudy (saifoo.khan):

\[\frac{(x-3)(x+3)}{x^2} \times \frac{x^2-3x}{(x-3) (x+4) }\]

OpenStudy (anonymous):



OpenStudy (anonymous):